OpenStudy (anonymous):

Find the percent of each number. 29% of 500.

OpenStudy (hoblos):

same way... 0.29*500=..

OpenStudy (anonymous):

But its confusing thoguh because some you add another 0 as for the 1.9 you did 0.019

OpenStudy (hoblos):

you divide it by 100 1.9/100 = 0.019 29/100 = 0.29